# Sub-diffraction estimation, discrimination and learning of quantum states of light

### v1

@@ IV.2 Multiple classes
In this Section we present results beyond binary classification. We consider MNIST objects from all classes of digits from 00 to $9$. To test the accuracy of SPADE for this multi-class problem we considered as features the moments of order two in both the Cartesian [see Eqs. (39)-(40)] and diagonal bases [Eqs. (41)-(42)]. In our Monte Carlo simulations, we assume that upon $N$ photon detected, half are measured in the Cartesian basis and half in the diagonal basis.
@@
The obtained confusion matrix is shown in Fig. 7 for $N=5000$ and $\tilde{\sigma}=10$, where each row is associated with the ground truth and the columns represent the guessed class. The matrix shows mixed results. As expected, SPADE is able to discriminate very well between 00’s and $1$’s. However, a lower accuracy is obtained for other digits. An example is given by the classes of $6$’s and $9$’s which are essentially indistinguishable from their second moments. This result is easily explained by noticing that, whereas the 00’s and $1$’s are well separated by the second moments, this is not in general the case for the other digits. This suggests that we may need to go beyond the second moments, i.e., include in our simulation and detection HG modes of higher degrees above $\text{HG}_{01}$ and $\text{HG}_{10}$.
IV.3 Higher moments
To extend the use of SPADE beyond binary classification, we come back to the expansion of the quantum states in the HG basis in Eq. (20):
$\rho$ $=\sum_{x,y}I_{x,y}e^{-\frac{|x|^{2}+|y|^{2}}{4\sigma^{2}}}$
×$\phantom{=}~{}\times\sum_{mm^{\prime}nn^{\prime}}\frac{x^{m+m^{\prime}}y^{n+n^{\prime}}}{(2\sigma)^{m+m^{\prime}+n+n^{\prime}}}\frac{|\text{HG}_{mn}\rangle\langle\text{HG}_{m^{\prime}n^{\prime}}|}{\sqrt{m!m^{\prime}!n!n^{\prime}!}}\,,$ (50)
and keep the terms up to the fourth order in $x/\sigma$ and $y/\sigma$.
The expansion includes basis vectors $|\text{HG}_{20}\rangle$, $|\text{HG}_{11}\rangle$, $|\text{HG}_{02}\rangle$. We obtain
$\rho$ $\simeq\left(1-\frac{\langle x^{2}\rangle+\langle y^{2}\rangle}{(2\sigma)^{2}}-\frac{\langle x^{4}\rangle+\langle x^{2}y^{2}\rangle+\langle y^{4}\rangle}{(2\sigma)^{4}}\right)|\text{HG}_{00}\rangle\langle\text{HG}_{00}|$ (51)
$+\frac{\langle x^{2}\rangle}{(2\sigma)^{2}}\left(|\text{HG}_{10}\rangle\langle\text{HG}_{10}|+\frac{|\text{HG}_{00}\rangle\langle\text{HG}_{20}|+\text{h.c.}}{\sqrt{2}}\right)+\frac{\langle y^{2}\rangle}{(2\sigma)^{2}}\left(|\text{HG}_{01}\rangle\langle\text{HG}_{01}|+\frac{|\text{HG}_{00}\rangle\langle\text{HG}_{02}|+\text{h.c.}}{\sqrt{2}}\right)$ (52)
[…]
$+\frac{\langle x^{4}\rangle}{(2\sigma)^{4}}\frac{|\text{HG}_{20}\rangle\langle\text{HG}_{20}|}{2}+\frac{\langle x^{3}y\rangle}{(2\sigma)^{4}}\frac{|\text{HG}_{11}\rangle\langle\text{HG}_{20}|+\text{h.c.}}{\sqrt{2}}$ (56)
$+\frac{\langle x^{2}y^{2}\rangle}{(2\sigma)^{4}}\left(|\text{HG}_{11}\rangle\langle\text{HG}_{11}|+\frac{|\text{HG}_{20}\rangle\langle\text{HG}_{02}|+\text{h.c.}}{2}\right)$ (57)
$+\frac{\langle xy^{3}\rangle}{(2\sigma)^{4}}\frac{|\text{HG}_{11}\rangle\langle\text{HG}_{02}|+\text{h.c.}}{\sqrt{2}}+\frac{\langle y^{4}\rangle}{(2\sigma)^{4}}\frac{|\text{HG}_{02}\rangle\langle\text{HG}_{02}|}{2}\,.$ (58)
@@ IV.3 Higher moments
By applying suitable measurements to the above state, one can obtain information about all the moments of the source intensity distribution up to the fourth order. As a concrete example, we focus on the problem of classifying MNIST objects belonging to the classes of $6$’s and $9$’s. Objects from these classes have essentially the same distribution of the second moments. Therefore, SPADE in the lower-order HG modes cannot classify them, as shown by the confusion matrix in Fig. 7. However, they are well separated by the third moments, as seen from Fig. 8. This suggests to extract information about the third moments $\langle x^{3}\rangle$ and $\langle y^{3}\rangle$ from the HG measurements. However, photon detection in the HG SPADE modes does not allow us to access the moments of the third order. We need to consider linear combinations of the HG basis elements. We consider the six vectors (this choice is not necessarily optimal, but provides a way to extract information about the third moments):
$\left\{|\text{HG}_{00}\rangle,\frac{|\text{HG}_{01}\rangle\pm|\text{HG}_{02}\rangle}{\sqrt{2}},\frac{|\text{HG}_{10}\rangle\pm|\text{HG}_{20}\rangle}{\sqrt{2}},|\text{HG}_{11}\rangle\right\}\,.$ (59)
Photon detection in these states is associated to the probabilities
$|\text{HG}_{00}\rangle\to 1-\frac{\langle x^{2}\rangle+\langle y^{2}\rangle}{(2\sigma)^{2}}-\frac{\langle x^{4}\rangle+\langle x^{2}y^{2}\rangle+\langle y^{4}\rangle}{(2\sigma)^{4}}\,,$ (60)
$\frac{|\text{HG}_{01}\rangle\pm|\text{HG}_{02}\rangle}{\sqrt{2}}\to\frac{1}{2}\left(\frac{\langle y^{2}\rangle}{(2\sigma)^{2}}+\frac{\langle y^{4}\rangle}{2(2\sigma)^{4}}\pm\frac{\sqrt{2}\langle y^{3}\rangle}{(2\sigma)^{3}}\right)\,,$ (61)
$\frac{|\text{HG}_{10}\rangle\pm|\text{HG}_{20}\rangle}{\sqrt{2}}\to\frac{1}{2}\left(\frac{\langle x^{2}\rangle}{(2\sigma)^{2}}+\frac{\langle x^{4}\rangle}{2(2\sigma)^{4}}\pm\frac{\sqrt{2}\langle x^{3}\rangle}{(2\sigma)^{3}}\right)\,,$ (62)
$|\text{HG}_{11}\rangle\to\frac{\langle x^{2}y^{2}\rangle}{(2\sigma)^{4}}\,.$ (63)
A partial optimisation could be obtained by introducing an angle of rotation:
@@
$(\sin{\varphi})^{2}\frac{\langle y^{2}\rangle}{(2\sigma)^{2}}+(\cos{\varphi})^{2}\frac{\langle y^{4}\rangle}{2(2\sigma)^{4}}+\sqrt{2}\sin{\varphi}\cos{\varphi}\frac{\langle y^{3}\rangle}{(2\sigma)^{3}}\,,$ (64)
@@
$(\cos{\varphi})^{2}\frac{\langle y^{2}\rangle}{(2\sigma)^{2}}+(\sin{\varphi})^{2}\frac{\langle y^{4}\rangle}{2(2\sigma)^{4}}-\sqrt{2}\sin{\varphi}\cos{\varphi}\frac{\langle y^{3}\rangle}{(2\sigma)^{3}}\,.$ (65)
The angle $\varphi$ can be optimised to maximise the signal-to-noise ratio, where the signal is represented by the moment of order three and the noise by the moments of order two and four. By putting
@@
$\sin{\varphi}|\text{HG}_{01}\rangle+\cos{\varphi}|\text{HG}_{02}\rangle\to\frac{y^{2}+y^{4}+2y^{3}}{(2\sigma)^{2}+2(2\sigma)^{4}}\,,$ (68)
$\cos{\varphi}|\text{HG}_{01}\rangle-\sin{\varphi}|\text{HG}_{02}\rangle\to\frac{2(2\sigma)^{2}y^{2}+\frac{y^{4}}{2(2\sigma)^{2}}-2y^{3}}{(2\sigma)^{2}+2(2\sigma)^{4}}\,.$ (69)
In our simulation, we use the above states, but we do not enforce the limit of small $x/\sigma$ and $y/\sigma$. This implies that exponential terms $e^{-\frac{-|x|^{2}-|y|^{2}}{4\sigma^{2}}}$ in (IV.3) are not replaced by $1$. Finally, to ensure a proper normalisation, we also include in the simulation the probabilities of detecting a photon in the higher-order terms
$\text{HG}_{03}$ $\to\frac{\langle y^{6}\rangle}{(2\sigma)^{6}}\frac{1}{6}\,,$ (70)
$\text{HG}_{12}$ $\to\frac{\langle x^{2}y^{4}\rangle}{(2\sigma)^{6}}\frac{1}{2}\,,$ (71)
$\text{HG}_{21}$ $\to\frac{\langle y^{2}x^{4}\rangle}{(2\sigma)^{6}}\frac{1}{2}\,,$ (72)
$\text{HG}_{30}$ $\to\frac{\langle x^{6}\rangle}{(2\sigma)^{6}}\frac{1}{6}\,.$ (73)
The results of our numerical experiments are shown in Fig. 9. First we have applied this modified SPADE using the above six features (with the optimised angle $\varphi$), with $N$ photon detection events, in a RF model. The accuracy is lower than in the case of 00’s vs $1$’s, and it requires many more detection events (higher values of $N$). This is because the third-order moments are suppressed by a factor proportional to $\tilde{\sigma}^{-3}$. With increasing $\tilde{\sigma}$, it is much less likely that is detected in an HG mode of higher degree, as most of them end up in the lower HG modes.
We then compare with DI, powered by FCNN. For small values of $\tilde{\sigma}$, DI works better than SPADE, but for larger values of $\tilde{\sigma}$ SPADE has better performance in terms of both accuracy and stability. Overall, this modified SPADE remains superior in the deep sub-Rayleigh regime (large values of $\tilde{\sigma}$), but more photon detection events are needed to acquire a decent signal-to-noise ratio.
@@ IV.4.1 MNIST Dataset
IV.4.2 Data Preprocessing and Models Architecture
For binary classification, starting from the full MNIST dataset we consider the reduced problem of binary classification: out of the $10$ classes, corresponding to the first ten integer numbers, we select only two classes. For the direct intensity case, the images are padded on an expanded grid of size $80\times 80$ pixels: this operation is necessary for preventing the combination of high scaling factor and blurring to cut out of the original grid useful information, which might lead to biased results. For the sake of computational time, which is greatly enhanced by the padded grid, we take a sub-set of the original dataset, reducing the size to $2000$ entries, balancing each of the two classes considered. Although this fact might, in principle, hinder the learning capabilities of the algorithm, both the trend of the loss function, and of the accuracy, are not effected, as they depend strongly upon the representation power of the feature vectors. For both DI and HG SPADE protocols we evaluate the performances for several manifolds of independent photon detection, ranging from low photo-detection frequency, i.e. $N=100$, to higher values of $N$. The upper values of $N$ selected, depend on whether or not using third order moments is necessary: in fact the probability of populating third order moments is low hence a greater collection of photons is necessary in order for these features to be meaningful for the classification. For SPADE, the feature vector is five-dimensional when moments up to the second order are considered, with entries extracted from the modes HG01, HG10, HG11, HG02, HG20. When third order moments are necessary, as in case of $6$ vs. $9$ classifcation, the SPADE feature vector is ten-dimensional, with the additional entries extracted from the modes, HG21, HG12, HG30, HG03, HG02, HG20, HG11.
@@ IV.4.2 Data Preprocessing and Models Architecture
IV.4.3 Training Procedure
The reduced MNIST dataset is divided into a training set and a validation set with an $70:30$ split. The training set is used to find the most suitable parameters for the model considered, while the validation set is employed to monitor the models performances. For FCNN, early stopping and dropout are in place, to prevent over-fitting. Finally, for the training procedure, the Adam optimiser has been used. The RF model minimises the loss function according to the Gini criterion Daniya et al. (2020), which measures the probability of extracting a sample out of the two classes. The FCNN instead is trained to minimise the binary cross entropy between the target and the output of the model:
@@ IV.4.3 Training Procedure
where $x_{n}$ is the target class and $y_{n}$ is the output class generated by the models. The hyperparamters selected for each model are listed in the result section.
@@ V Conclusions
The Rayleigh resolution criterion heuristically states that, through imaging, it is hard to resolve details below the Rayleigh length, which is determined by the wavelength and by the size of the pupil of the optical imaging system. However, this criterion only applies to imaging systems based on direct imaging (DI), where one measures the intensity of the focused field on the image plane. By contrast, interferometric measurements can be used to obtain information not only about the intensity but also about the phase of the field. Measurement techniques based on spatial-mode demultiplexing (SPADE) have been recently investigated and demonstrated to achieve sub-Rayleigh resolution in the context of image discrimination and parameter estimation Tsang et al. (2016); Pushkina et al. (2021); Řehaček et al. (2017); Zhou and Jiang (2019); Yu and Prasad (2018); Paúr et al. (2018); Tham et al. (2017); Santamaria et al. (2024); Paúr et al. (2016); Rouvière et al. (2024).
In this work, we have explored the use of SPADE as a physical pre-processing of the quantum field, where first the field is measurement using SPADE, and then the collected data are fed into a machine learning algorithm. Our approach can be interpreted as a hybrid quantum-classical machine learning algorithm, where the physical part consists in hardware pre-processing of the optical field, followed by classical post-processing through algorithms such as Random Forest (RF) or Fully Connected Neural Network (FCNN). We have applied this approach to classify objects from the MNIST database of handwritten digits from 00 to $9$. Our simulations and numerical experiments show that SPADE can classify objects that are otherwise indistinguishable using classical approaches based on DI.
Our theoretical analysis elucidates the special role played by the Hermite-Gaussian (HG) modes of the transverse field. Source objects that are well-separated by the second moments of their intensity distribution are classified with high accuracy by photon detection in the lower HG modes. The same applies to all moments of even orders. However, the probability of detecting a photon in higher-order HG modes is exponentially suppressed in the sub-Rayleigh regime. For moments of odd order, we need to consider a modified SPADE obtained by linear superposition of the HG modes. As an example, we have considered the case of $6$ vs $9$ classification, as these two classes are well-separated by their third moments.
@@
To further improve the accuracy of our hybrid image classifier, one should optimise the choice of the demultiplexer that is used to implement SPADE. Formally, this means optimise the choice of the unitary matrix that mathematically represents such a transformation. That is, during the training phase, not only the parameters of the neural network are optimised, but also those of the unitary matrix that represents the multiplexer.
@@
Effects of experimental imperfections on the efficacy of SPADE, in particular cross-talk, have been assessed by a number of previous works, e.g. Gessner et al. (2020); Boucher et al. (2020); Santamaria et al. (2023); Len et al. (2020); Lupo (2020); Oh et al. (2021). In general noise and cross-talk decrease the efficacy of SPADE, but the advantage with respect to DI remains for mild imperfections. We expect the same to happen in our machine learning framework, but a detailed analysis is left for future works. Finally, we have focused on supervised learning but our approach could be also extended to un-supervised learning, as well as to classes of objects of higher complexity than the MNIST digits.
Acknowledgements.

### v2

@@ IV.2 Multiple classes
Here we consider MNIST objects from all classes of digits from 00 to $9$. To test the accuracy of SPADE for this multi-class problem we considered features arising from a measurement of lower HG modes, in both the HG basis and its rotated version.
For the rotated HG basis, the probabilities are given in Eq. (28)-(29). For the un-rotated basis, photon detection is simulated in the lower-order HG modes:
$\left\{|\text{HG}_{00}\rangle,|\text{HG}_{10}\rangle,|\text{HG}_{01}\rangle\right\}\,,$ (34)
yielding the probabilities $p_{00}^{\text{SPADE}}$ in (28) and
$|\text{HG}_{10}\rangle$ $\to\frac{1}{4\sigma^{2}}\sum_{x,y}I_{x,y}x^{2}e^{-\frac{|x|^{2}+|y|^{2}}{4\sigma^{2}}}=p_{10}^{\text{SPADE}}\,,$ (35)
$|\text{HG}_{01}\rangle$ $\to\frac{1}{4\sigma^{2}}\sum_{x,y}I_{x,y}y^{2}e^{-\frac{|x|^{2}+|y|^{2}}{4\sigma^{2}}}=p_{01}^{\text{SPADE}}\,.$ (36)
In our Monte Carlo simulations, we assume that upon $N$ photons detected, half are measured in the Cartesian basis and half in the diagonal basis.
@@
The obtained confusion matrix is shown in Fig. 5 for $N=5000$ and $\tilde{\sigma}=10$, where each row is associated with the ground truth and the columns represent the guessed class. The matrix shows mixed results. As expected, SPADE is able to discriminate very well between 00’s and $1$’s. However, a lower accuracy is obtained for other digits. An example is given by the classes of $6$’s and $9$’s, which are essentially indistinguishable. This result is easily explained by noticing that, whereas the 00’s and $1$’s are well separated by the second moments, this is not in general the case for the other digits. This suggests that we may need to go beyond the second moments, i.e., include in our simulation and detection HG modes of higher degrees above $\text{HG}_{01}$ and $\text{HG}_{10}$.
IV.3 Higher-order HG modes
To extend the use of SPADE beyond second moments, we include in our simulations the events of photon detection in higher-order HG vectors, obtained from linear combinations of $|\text{HG}_{20}\rangle$, $|\text{HG}_{02}\rangle$.
@@ IV.3 Higher-order HG modes
As a first concrete example, we focus on the problem of classifying MNIST objects belonging to the classes of $6$’s and $9$’s. Objects from these classes have essentially identical distributions of the second moments, but are well separated by the third moment $\langle y^{3}\rangle$, as seen from Fig. 6. We therefore consider measurements in the six vectors (this choice is not necessarily optimal, but provides a way to extract information about $\langle y^{3}\rangle$):
$\left\{|\text{HG}_{00}\rangle,\frac{|\text{HG}_{01}\rangle\pm|\text{HG}_{02}\rangle}{\sqrt{2}},|\text{HG}_{10}\rangle,|\text{HG}_{20}\rangle,|\text{HG}_{11}\rangle\right\}\,.$ (37)
Photon detection in these states is related to combination of moments up to the fourth order, including
$\frac{|\text{HG}_{01}\rangle\pm|\text{HG}_{02}\rangle}{\sqrt{2}}$
$\to\frac{1}{2}\sum_{x,y}I_{x,y}e^{-\frac{|x|^{2}+|y|^{2}}{4\sigma^{2}}}\left(\frac{y^{2}}{(2\sigma)^{2}}+\frac{y^{4}}{2(2\sigma)^{4}}\pm\frac{\sqrt{2}y^{3}}{(2\sigma)^{3}}\right)$ (38)
$\sim\frac{1}{2}\left(\frac{\langle y^{2}\rangle}{(2\sigma)^{2}}+\frac{\langle y^{4}\rangle}{2(2\sigma)^{4}}\pm\frac{\sqrt{2}\langle y^{3}\rangle}{(2\sigma)^{3}}\right)\,.$ (39)
A partial optimisation may be obtained by introducing an angle of rotation:
@@
$\hskip 8.5359pt(\sin{\varphi})^{2}\frac{\langle y^{2}\rangle}{(2\sigma)^{2}}+(\cos{\varphi})^{2}\frac{\langle y^{4}\rangle}{2(2\sigma)^{4}}+\sqrt{2}\sin{\varphi}\cos{\varphi}\frac{\langle y^{3}\rangle}{(2\sigma)^{3}}\,,$ (40)
@@
$\hskip 8.5359pt(\cos{\varphi})^{2}\frac{\langle y^{2}\rangle}{(2\sigma)^{2}}+(\sin{\varphi})^{2}\frac{\langle y^{4}\rangle}{2(2\sigma)^{4}}-\sqrt{2}\sin{\varphi}\cos{\varphi}\frac{\langle y^{3}\rangle}{(2\sigma)^{3}}\,.$ (41)
The angle $\varphi$ is optimised to maximise the signal-to-noise ratio, where the signal is represented by the moment of order three and the noise by the moments of order two and four. By putting
@@
$\sin{\varphi}|\text{HG}_{01}\rangle+\cos{\varphi}|\text{HG}_{02}\rangle$ $\to\frac{\langle y^{2}\rangle+\langle y^{4}\rangle+2\langle y^{3}\rangle}{(2\sigma)^{2}+2(2\sigma)^{4}}\,,$ (44)
$\hskip 2.84544pt\frac{2(2\sigma)^{2}\langle y^{2}\rangle+\frac{\langle y^{4}\rangle}{2(2\sigma)^{2}}-2\langle y^{3}\rangle}{(2\sigma)^{2}+2(2\sigma)^{4}}\,.$ (45)
Finally, to ensure normalisation, we also include in the simulation the probabilities of detecting a photon in the higher-order terms $\text{HG}_{03}$, $\text{HG}_{12}$, $\text{HG}_{21}$, $\text{HG}_{30}$.
The results of our numerical experiments are shown in Fig. 7. We have applied this modified SPADE using the above features (with the optimised angle $\varphi$), with $N$ photon detection events, in a RF model. The accuracy is lower than in the case of 00’s vs $1$’s, and it requires many more detection events (higher values of $N$). This is because the third-order moments are suppressed by a factor proportional to $\tilde{\sigma}^{-3}$. With increasing $\tilde{\sigma}$, it is much less likely that a photon is detected in an HG mode of higher order, as most of them end up in the low-order HG modes.
We then compare with DI, powered by FCNN. For small values of $\tilde{\sigma}$, DI works better than SPADE, but for larger values of $\tilde{\sigma}$ SPADE has better performance in terms of both accuracy and stability. Overall, this modified SPADE remains superior in the deep sub-Rayleigh regime (larger values of $\tilde{\sigma}$), but more photon detection events are needed to acquire a decent signal-to-noise ratio. In conclusion, photons detected in the rotated HG modes are rare yet highly informative.
@@
Going beyond binary classification, we repeated the multi-class numerical experiment but this time using the modified SPADE. The results are shown in Fig. (8). While in general the performance depends on which specific class is best characterised by the second of third moments, the confusion matrix in Fig. (8) indicates that in most cases considering the higher-order HG modes yields higher accuracy in the multi-class classification problem.
Figure 8: Confusion matrix of the output of the RF for multi-class classification, performed on modified SPADE in Eq. (37). The figure refers to an effective blurring factor $\tilde{\sigma}=10$ and number of photon detection events $N=5000$. Compared with the confusion matrix in Fig. (5), this shows that including higher-order HG modes may substantially increase multi-class classification.
@@ IV.4.1 MNIST Dataset:
IV.4.2 Data Preprocessing and Models Architecture:
For binary classification, starting from the full MNIST dataset we consider the reduced problem of binary classification: out of the $10$ classes, corresponding to the first ten integer numbers, we select only two classes. For the direct intensity case, the images are padded on an expanded grid of size $80\times 80$ pixels: this operation is necessary for preventing the combination of high scaling factor and blurring to cut out of the original grid useful information, which might lead to biased results. For the sake of computational time, which is greatly enhanced by the padded grid, we take a sub-set of the original dataset, reducing the size to $2000$ entries, balancing each of the two classes considered. Although this fact might, in principle, hinder the learning capabilities of the algorithm, both the trend of the loss function, and of the accuracy, are not effected, as they depend strongly upon the representation power of the feature vectors. For both DI and SPADE protocols we evaluate the performances for several manifolds of independent photon detection, ranging from low photo-detection frequency, i.e. $N=100$, to higher values of $N$. The upper values of $N$ selected depend on whether or not using third order moments is necessary: in fact the probability of populating third order moments is low hence a greater collection of photons is necessary in order for these features to be meaningful for the classification. For SPADE, the feature vector is five-dimensional when moments up to the second order are considered, with entries extracted from the modes HG01, HG10, HG11, HG02, HG20. When third order moments are necessary, as in case of $6$ vs. $9$ classification, the SPADE feature vector is ten-dimensional, with the additional entries extracted from the modes, HG21, HG12, HG30, HG03, HG02, HG20, HG11.
@@ IV.4.2 Data Preprocessing and Models Architecture:
IV.4.3 Training Procedure:
The reduced MNIST dataset is divided into a training set and a validation set with an $70:30$ split. The training set is used to find the most suitable parameters for the model considered, while the validation set is employed to monitor the models performances. For FCNN, early stopping and dropout are in place, to prevent over-fitting. Finally, for the training procedure, the Adam optimiser has been used. The RF model minimises the loss function according to the Gini criterion Ginic , which measures the probability of extracting a sample out of the two classes. The FCNN instead is trained to minimise the binary cross entropy between the target and the output of the model:
@@ IV.4.3 Training Procedure:
where $x_{n}$ is the target class and $y_{n}$ is the output class generated by the models. The hyperparameters selected for each model are listed in the result section.
@@ V Conclusions
We have explored a hybrid scheme for image classification, where the physical part is a hardware-based pre-processing of the optical field, followed by photon-detection. The computational part is a machine learning algorithm, such as Random Forest (RF) or Fully Connected Neural Network (FCNN). The hardware part is represented by a multi-mode interferometer that implements spatial-mode demultiplexing (SPADE) of the transverse field. In particular, we have focused on Hermite-Gaussian (HG) modes and their linear combinations.
Our work is the first attempt to exploit SPADE in the context of image classification and machine learning within the sub-Rayleigh and photon-counting regime. With some exceptions, see e.g. Refs. Lvovsky , Frank:23 , most of the previous works were limited to problems of parameter estimation and hypothesis testing.
As a case study, we have applied this approach to classify objects from the MNIST database of handwritten digits from 00 to $9$. Our simulations and numerical experiments show that SPADE can classify objects that are otherwise indistinguishable using classical approaches based on DI.
@@
To further improve the accuracy of our hybrid image classifier, one should optimise the choice of the demultiplexer that implements SPADE. Formally, this means optimise the choice of the unitary matrix that mathematically represents such a transformation. That is, during the training phase, not only the parameters of the neural network are optimised, but also those of the unitary matrix Spall .
@@
Effects of experimental imperfections on the efficacy of SPADE, in particular cross-talk, have been assessed by a number of previous works, see for example Refs. PhysRevLett.125.100501 , Boucher2020 , Santamaria22 , Banaszek2020 , PhysRevA.101.022323 , PhysRevLett.126.120502 . In general noise and cross-talk decrease the efficacy of SPADE, but the advantage with respect to DI remains for mild imperfections. We expect the same to happen in our machine learning framework, but a detailed analysis is left for future works. We note that our approach may be combined with other schemes of quantum imaging and stellar interferometry, e.g. app11146414 , PhysRevLett.109.070503 , PhysRevLett.123.070504 , PhysRevLett.129.210502 . Finally, our approach could be also extended to un-supervised learning, as well as to classes of objects of higher complexity than the MNIST digits.
VI Acknowledgements